Page tree

Versions Compared

Key

  • This line was added.
  • This line was removed.
  • Formatting was changed.

...

LTP6 Block STP cycleIOR delivery opensIOR deadlineSTP POR deadline (on SOC) Estimated/Planned uplink by MOC (MIBsStart of STP cycle onboardComments

SOC ESA responsible

Main operator
1183

 

 

 


MIB from 18/19  Nov

 

MAINT (SA_REL)

Synoptics

Operating: 

All Instruments except PHI.

SPICE cal, flat and dark. STIX analysis.

Anik De Groof 

usual SCO support

Javier Arenas 

1184

 

 

 

 

MIB from 18/19  Nov

 

Synoptics

Operating: 

All Instruments except PHI (flush only)

SPICE cal, flat. Testing in parallel to WOL on first day

STIX analysis

METIS HSD CLOSING

(warning) Note: all instruments should switch off before 9-Jan pass for SSMM resizing, except:

    • EUI, SWA, MAG can stay SAFE
    • SPICE can stay STANDBY with the particle-deflector system enabled until the HSD closes (in next STP)
David Williams 

usual SCO support

Fernando Martin Porqueras 

1185

 

 

1200utc   => feedback on PORs from MOC EOB

 (Fri) or weekend?

MIB from 18/19  Nov

 

(warning) NEAR_SUN conditions start 15/1/22

EUI (FSI), SPICE, PHI (HRT/FDT), HSD CLOSE

(warning) MAINT/ENG (instruments off, except EUI, MAG and SWA in safe)

EUI(FSI), SPICE, METIS PHI (HRT/FDT), HSD OPEN

Synoptics

Operating: All Instruments except PHI.

SPICE cal, flat and dark. STIX analysis.

usual SCO support until 23 Dec. DM required, STP issues reviewed as part of DM+

Fernando Martin Porqueras 

1186

 

 

 

  or , in case of new MIB for 189, earlier date preferred.

MIB from 18/19  Nov

 

TCM type 2

MAINT (any) in NEAR SUN conditions

All instruments operating (RS synoptics)

PHI helioseismology campaign (= PHI_RAW_FDT_6)

spice cal, flat, dark. 

David Williams and Jayne Lefort  after Dave is on leave


on call support 23 dec -6 jan.

from 7 Jan (inc)  usual SCO support

Javier Arenas 

1187

 

 

 

 

TBC Target MIB for 187 and block 2 (after 14 Jan), ideally possible to apply at SOC from STP 188.

 

SA_ROT

All instruments operating (RS synoptics)

PHI processing of helioseismology campaign data

spice cal, flat, dark. 


Anik De Groof 

from 7 jan usual sco support

Diego Fraga Agudo / Mar Sierra 

2188

/  

 

 

 


 

EUI HRI HSD opening at end STP

All instruments operating (RS synoptics)

(warning) Metis test after switch-over to new FSW

spice cal, flat, dark. 

PHI raw synoptics run.

2189

 

 

 


 

SA ROT

All instruments operating (RS synoptics)

Jitter test during WOL, incl SPICE, PHI and EUI

(warning) EUI runs validation campaign for 6-8 March campaigns

SPICE HSD CLOSING ahead of LUT upload in next STP

2190

 

 

 

TBC Target MIB update for 190 and block 3  (ideally to apply at SOC 191 if possible)

 

TCM type 2

MAINT (any) in NEAR SUN conditions

SA Rotation

all instruments on (RS synoptics)

SPICE HSD opening after LUT update

SPICE CE heater set to op limits (i.e., goes cold)

spice cal, flat, dark. 

Jayne Lefort 
31913/4 February 

 

 

 


 

All payload on and operating in RS synoptics or IS science modes.

2 limb-pointing campaigns to East and later West limb for coordinated science with PSP

RSW_EXT: SPICE/PHI/MET/EUI can plan for (=reserve window in STP IOR) and send VSTP IORs if needed

31923/4 February 

 

(wed)

(mon)  or weekend?


  06:00

(warning) Long STP (~12d)

RSW_EXT and start RSW1: VSTP IORs can be planned for and sent, during RSWs only

All Payload in science (and calibration) modes

Coordinated science activities (SOOPs):

  • Connection mosaic incl offpointing to 3 positions (pVSTP)
  • Slow wind Connection incl target tracking (pVSTP)
  • Polar observations incl pole pointing (pVSTP)
  • Nanoflare observations at active target (pVSTP)
  • Full disk mosaic at 25 pointing positions (pVSTP)
  • Coronal He abundance in coord with sounding rocket (sun centre)
  • Nanoflare observations at sun centre
  • Bright points SOOP at sun centre

(warning) Metis HSD closure and instrument safing during off-pointing within 0.52 AU 

Calibration activities:

  • EUI/PHI flatfield (3/3/22) incl 8 offpointing positions
  • EUI/SPICE cross calibration (7/3/22)
  • PHI FDT and HRT & SPICE calibrations

Synoptics at end RSW.

(note inferior conjunction does not impose operational constraints)


(warning) SPICED Heater switching off, HIS needs to be switched OFF for this.  13th March during WOL

31933/4 February 

  (tue)

  (mon)

  (thur)

TBC Target MIB for 193, and  block 4 ideally to apply at SOC 194 if possible 


(warning) Short STP (<5d)

MAINT ANY + TCM Type-2 (in near-sun conditions)

SA rotation

(warning) SWA_TCS_NEAR

RSW_EXT: SPICE/PHI/MET/EUI can plan for (=reserve window in STP IOR) and send VSTP IORs if needed

SPICE HSD operations for LUT update
Metis HSD closure for incompatible SC pointing during flatfield and start RSW2 + dark calibration

Calibration activities between RSW1 & 2:

  • CALIB_ROLL_RS incl 8 roll angles
  • CALIB_OFFPOI_STRAYLIGHT incl 22 offpointing positions
  • EUI/PHI flatfield incl 8 offpointing positions
  • Metis dark calibation
Anik De Groof 
419428 Feb (latest)

(thur)

(wed)

(mon)


16:30

RSW_EXT and Start RSW2+3: VSTP IORs can be planned for and sent, during RSWs

Frequent WOLs.

All payload On and operating except Metis (warning)

Coordinated science activities (SOOPs):

  • Nanoflare observations at Quiet and Active target (pVSTP)
  • Slow wind Connection incl target tracking (pVSTP)


419528 Feb (latest)

(fri) or weekend?


 

RSW2+3: VSTP IORs can be planned for and sent, during whole STP.

(warning) Frequent WOLs, interrupting science operations for several instruments.

All payload On and operating.

Metis HSD opening for
coordinated science campaigns (SOOPs), at disk centre:

  • 4 instances of Coronal Dynamics SOOP
  • Eruption Watch SOOP
  • Coronal Hole boundary expansion SOOP


419628 Feb (latest)

 

 

(fri) or weekend?


 

RSW2+3: VSTP IORs can be planned for and sent, during whole STP.

Frequent WOLs.

All payload On and operating.

(warning) Metis HSD closure and instrument safing during most of the STP.

Coordinated science activities (SOOPs):

  • Eruption Watch, incl Metis, at disk centre
  • Nanoflare observations at Active target (pVSTP)
  • Polar Observations incl pole pointing (pVSTP)
  • Connection Mosaic incl 6 offpointing positions (pVSTP)
  • ~4 days long AR tracking (pVSTP)
  • Polar Observations incl pole pointing & disk centre pointing for Metis context (pVSTP) 


4419728 Feb (latest)

(thur)

(wed)

(mon)

Target MIB for next block? if needed available   ideally applicable MOC STP 197 and  STP SOC 198

 

(warning) SWA_TCS_FAR

EUI/HRI HS door closure after end RSW3.

All payload On and operating.

PHI/EUI flatfield calibration, incl 8 offpointing positions

Synoptics.

SWA must be off at end of STP in preparation for the WOL at start of next LTP

The SPICE  heater should be powered on during the WOL of 10-Apr-22 12:05–15:05, with the SOU heater being switched back on at 12:35 His off